OpenStudy (anonymous):

Help! For the function y = f(x), what is the ordered pair for the point on the graph when x = b - 2? (b - 2, f(b -2)) (x, f(b)) (x, b - 2) (b - 2, f(b))

OpenStudy (anonymous):

@cample_st

OpenStudy (anonymous):

@camble_st

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

will in y = f(x) replace x with b - 2

OpenStudy (campbell_st):

so x = b - 2 y = f(b -2)

OpenStudy (anonymous):

So it's A

OpenStudy (campbell_st):

yes